# Assessing Physics Students’ Scientific Argumentation using Natural Language Processing

### v1

@@ 6 Conclusions
To address our first research question: by introducing more levels of scaffolding throughout the semesters, we found that students constructed arguments focused more on physics concepts opposed to mathematical steps. To address our second research question: we were able to use unsupervised NLP methods to assess student argumentation in the context of physics problem-solving.
By investigating the top words of the topics and the counts of essays in each topic, we were able to see a change in student focus across four semesters. In the semesters, with less or no scaffolding students seemed to focus equally across topics, with some topics focusing on superficial features, mathematical concepts, or physics concepts. In the latest two semesters, we saw more essays focus on energy principle concepts than any other topic in the respective semesters. In these semesters, students received the same levels of scaffolding, which was more than the previous two. This supports that we were able to improve student argumentation and track the progress over the course of two years. In addition, we measured the coherence factors and see that later semesters have more coherence that that of Fall 2022.
6.1 Implications and Future Work
The results of this study show promise for exploring various unsupervised machine learning approaches to compare students across different semesters. We intend to expand this work to track student scientific arguments throughout a single semester to determine if we can capture their argumentation progress using unsupervised machine learning techniques.
@@ 7 Acknowledgments
This research is supported in part by U.S. National Science Foundation grants 2111138 and 2300645 Any opinions, results, and findings expressed here are those of the authors and not of the Foundation.

### v2

@@ 6 Conclusions
In summary, using ML, we analyzed the top words, distribution of topics, and the concentration factor of the topics obtained from NMF. Through this analysis and across all student responses, we were able to identify topics that reflected on energy concepts, mathematical steps, surface features, etc. that all matched the context of the given problem. To address our research question: How does the quality of students’ scientific argumentation in the context of physics problem-solving evolve in response to scaffolding and fading over multiple weeks of instruction? we examined the top words and distribution of topics over four semesters of varying argumentation scaffolds to detect any changes. In semesters with minimal scaffolding (Fall 2022 and Spring 2023), students’ arguments were distributed more evenly across topics, indicating an equal emphasis on both conceptual and procedural elements. In contrast, in semesters with greater scaffolding (Fall 2023 and Spring 2024), topics were more concentrated around conceptually rich ideas, particularly those emphasizing energy principles.These results suggest that explicit argumentation scaffolds can help students shift from procedural reasoning toward more conceptually grounded argumentation in physics.
By investigating both the top words and topic distribution across semesters, we observed a clear change in student focus as instructional scaffolds were implemented. In the earlier semesters, student focus was distributed equally across conceptual and procedural topics. In the later semesters, we observed their focus to concentrate around physics principles, such as conservation of energy, associated with the problem statement. Overall, these findings indicate that structured scaffolding in scientific argumentation instruction supports students in constructing more conceptually meaningful explanations over time. These findings demonstrate that unsupervised ML methods, such as topic modeling, can capture large-scale trends in student argumentation, offering a scalable and reproducible approach for characterizing conceptual versus procedural focus in written physics responses.
6.1 Limitations, Implications, and Future Work
The results of this study highlight the potential of unsupervised machine learning methods as tools for evaluating student written argumentation at scale. While promising, this approach is limited by the interpretive nature of topic modeling and the need for human validation to ensure meaningful topic labeling. Nevertheless, these methods can serve as efficient diagnostic tools for instructors seeking to track changes in student reasoning across semesters. There were also inherent limitations related to the implementation of scaffolds within the course. First, scientific argumentation was taught and scaffolded only within the recitation section of the course; students did not receive any formal instruction on scientific argumentation from lecture or lab. This restricted the extent of student exposure and may have limited opportunities for reinforcement. Second, since scaffolds occurred exclusively in recitation, the quality and depth of instruction was partly reliant upon GTA discretion. Although GTAs received weekly training and were provided with standardized slides introducing scientific argumentation, the extent to which they discussed these concepts varied. Nevertheless, all students received identical written scaffolds through the recitation worksheets, which helped maintain a baseline level of consistency in the instructional material.
Future work will expand this analysis to track individual students’ scientific arguments longitudinally within a single semester to explore whether similar trends in conceptual focus can be detected over shorter timescales. We also plan to extend the dataset to include additional semesters to examine whether the observed differences between fall and spring cohorts persist. More broadly, this line of work points toward the development of machine learning–based frameworks that can provide instructors with scalable, evidence-based feedback on student scientific argumentation in physics and beyond.
@@ 7 Acknowledgments
This research is supported in part by U.S. National Science Foundation grants 2111138. Any opinions, results, and findings expressed here are those of the authors and not of the Foundation.